# ColNet: Collaborative Optimization in Decentralized Federated Multi-task Learning Systems

### v1

@@ Abstract
The integration of Federated Learning (FL) and Multi-Task Learning (MTL) has been explored to address client heterogeneity, with Federated Multi-Task Learning (FMTL) treating each client as a distinct task. However, most existing research focuses on data heterogeneity (e.g., addressing non-IID data) rather than task heterogeneity, where clients solve fundamentally different tasks. Additionally, much of the work relies on centralized settings with a server managing the federation, leaving the more challenging domain of decentralized FMTL largely unexplored. Thus, this work bridges this gap by proposing ColNet, a framework designed for heterogeneous tasks in decentralized federated environments. ColNet divides models into the backbone and task-specific layers, forming groups of similar clients, with group leaders performing conflict-averse cross-group aggregation. A pool of experiments with different federations demonstrated ColNet outperforms the compared aggregation schemes in decentralized settings with label and task heterogeneity scenarios.
1 Introduction
@@
The integration of Federated Learning (FL) and Multi-Task Learning (MTL) has gained significant attention for its ability to tackle diverse but related tasks in a collaborative and privacy-preserving manner. FL trains models across clients by exchanging model updates rather than raw data, preserving data privacy Kairouz et al. (2021). Decentralized FL (DFL) eliminates the need for a central server, enhancing robustness by avoiding single points of failure Beltrán et al. (2023). MTL, traditionally performed on a single machine, leverages task interdependence to improve generalization, reduce overfitting, and address data sparsity Caruana (1997); Zhang and Yang (2021). Extending MTL to federated settings results in Federated Multi-Task Learning (FMTL) Smith et al. (2017), combining FL’s privacy preservation with MTL’s collaborative task-sharing benefits.
@@
In both centralized and decentralized FMTL, existing research primarily addresses the challenge of non-independent and identically distributed (non-IID) data across clients, a core issue in FL Zhao et al. (2018). Non-IID data arises in various forms, including attribute skew (differences in feature distributions), label skew (differences in label distributions), and temporal skew (time-based changes in data distributions), each complicating model training Zhu et al. (2021). However, while much progress has been made in handling non-IID data, most studies focus on label heterogeneity rather than deeper task heterogeneity. Many approaches assume that all clients access the same set of class labels, even in non-IID settings. In reality, clients may encounter only subsets of labels relevant to their tasks Aoki et al. (2022), creating disparities in outputs and limiting collaborative learning. Addressing task heterogeneity, where clients solve fundamentally different problems, remains a significant challenge, especially in decentralized settings.
@@
Aggregation approaches like FedAvg struggle in multi-task learning scenarios due to their reliance on model homogeneity Huang et al. (2022). Task heterogeneity introduces complexities, as clients may require distinct model architectures for tasks (e.g., classification versus point detection), leading to conflicting and dominant gradients during training Lu et al. (2024); Yu et al. (2024). While recent methods such as FedBone Chen et al. (2023) and FedHCA2 Lu et al. (2024) address these challenges, they focus on centralized settings (CFMTL), leaving decentralized FMTL (DFMTL) largely unexplored. Decentralization amplifies these challenges due to the absence of a coordinating server, demanding robust methods for managing tasks and model diversity.
This work introduces ColNet, a decentralized framework designed to train multi-task models in heterogeneous environments. ColNet groups clients by tasks, enabling intra-group federated averaging of backbones layers. Task group leaders then exchange averaged backbones across groups using a hyper-conflict-averse (HCA) aggregation scheme to reduce conflicting gradients and enhance cross-task collaboration.
Table 1: Overview of Research in Federated Multi-Task Learning
@@
Convex MOCHA Smith et al. (2017) Convex optimization, dynamic client inclusion Label heterogeneity
Optimization OFMTL Li et al. (2019) Convex optimization, dynamic client inclusion Label heterogeneity
Regularization Ditto Li et al. (2021) Bi-level optimization, local regularization Label heterogeneity
PMTL Hu et al. (2021) Differential privacy, noise-injected global aggregation Label heterogeneity
Privatization FedPer Arivazhagan et al. (2019) Shared base layers, private personalization layers Label heterogeneity
Centralized FedRep Collins et al. (2021) Shared base layers, head freezing Label heterogeneity
Grouping CFL Sattler et al. (2020) Client grouping, cosine similarity Label heterogeneity
HeurFedAMP Huang et al. (2021) Dynamic grouping, cosine similarity Label heterogeneity
Knowledge EFDLS Xing et al. (2022) Knowledge distillation, local teacher-student models Label heterogeneity
Distillation FedICT Wu et al. (2023) Two-way knowledge distillation, reduced model updates Label heterogeneity
Hybrid FedBone Chen et al. (2023) Split learning, gradient projection, task adaptation Task heterogeneity
FedHCA2 Lu et al. (2024) Hyper conflict-averse aggregation, encoder-decoder architecture Task heterogeneity
Regularization DCLPMN Vanhaesebrouck et al. (2016) Graph-based regularization, neighbor collaboration Label heterogeneity
Decentralized Hybrid SpreadGNN He et al. (2022) Decentralized aggregation, molecular data focus Label heterogeneity
Hybrid ColNet (This work) Shared backbone, grouping, hyper conflict-averse aggregation Task heterogeneity
Experiments on CIFAR-10 and CelebA datasets demonstrate ColNet’s effectiveness. First, the framework’s performance is analyzed with varying aggregation round frequencies, showing that more frequent aggregations improve model performance. Second, different levels of layer privatization are examined, revealing that sharing most layers benefits class label heterogeneity setups, while increasing privatization slightly enhances performance in task heterogeneity scenarios. Finally, compared with other schemes, ColNet consistently outperforms alternatives in both class label and task heterogeneity scenarios, highlighting its capability to address diverse challenges in DFL.
2 Related Work
This section provides an overview of FMTL research, covering both centralized and decentralized approaches and comparing them based on the techniques employed and types of heterogeneity addressed, as presented in Table 1.
@@ A. Centralized Approaches
FMTL was first introduced in 2017 with the MOCHA framework Smith et al. (2017), and since then, a variety of approaches have emerged to tackle different aspects of multi-task learning in centralized federated settings. Early work primarily focused on solving convex optimization problems, such as MOCHA and OFMTL Li et al. (2019), which were robust but unsuitable for modern non-convex deep learning methods. Subsequent research addressed this limitation, extending CFMTL to handle non-convex problems and introducing techniques like regularization, privatization, grouping, and knowledge distillation.
Ditto Li et al. (2021) introduced bi-level optimization techniques that allow each client to maintain a personalized model while regularizing it toward a global model. This approach ensures uniform performance across devices and enhances resistance to data and model poisoning attacks. Similarly, PMTL Hu et al. (2021) leverages relaxed differential privacy to learn personalized models while safeguarding client data. By adding noise during global aggregation, PMTL enables effective collaboration without compromising privacy. FedPer Arivazhagan et al. (2019) tackled statistical heterogeneity by dividing models into shared base layers and private personalization layers. Extensions like FedRep Collins et al. (2021) introduced mechanisms such as freezing the shared layers during early training phases to improve adaptability and achieve higher test accuracy.
Frameworks like CFL Sattler et al. (2020) introduced grouping strategies to cluster clients with similar data distributions, reducing the impact of conflicting gradients during aggregation. HeurFedAMP Huang et al. (2021) further refined this concept by using cosine similarity to dynamically determine grouping, allowing more granular adaptation to client data heterogeneity. EFDLS Xing et al. (2022) utilized knowledge distillation models to transfer knowledge efficiently within individual clients. This method enables collaborative learning by sharing only essential updates, reducing communication overhead. FedICT Wu et al. (2023) extended this idea by employing a two-way distillation process that customizes local models for specific tasks while maintaining a flexible and adaptive global model, making it suitable for diverse client environments.
Most existing studies primarily focus on class label heterogeneity, restricting their adaptability in dynamic environments. Frameworks like FedBone Chen et al. (2023) introduced split learning to separate general models on a central server from task-specific models on clients. Similarly, FedHCA2 Lu et al. (2024) proposed innovative aggregation techniques, such as HCA Aggregation, to address gradient conflicts, but it was designed for centralized environments.
@@ B. Decentralized Approaches
DFMTL eliminates the need for a central server, allowing clients to communicate directly in a peer-to-peer manner. Despite the advantages, DFMTL introduces added challenges, such as managing gradient conflicts, asynchronous updates, and efficient client communication. Early research, such as DCLPMN Vanhaesebrouck et al. (2016), addressed these challenges by leveraging graph-based regularization to model task relationships, enabling collaboration between neighboring clients. SpreadGNN He et al. (2022) applied decentralized aggregation techniques tailored for tasks like molecular data analysis to improve scalability and address domain-specific requirements.
As presented in Table 1, while research on FMTL has made significant progress, most works addresses label heterogeneity and non-IID data rather than task heterogeneity, where clients solve fundamentally different tasks. Additionally, decentralized settings remain underexplored compared to centralized frameworks, despite their potential for improved privacy and scalability. Existing frameworks rarely address task heterogeneity in decentralized environments, leaving a clear gap for Heterogeneous DFMTL. This work addresses this gap by proposing a novel framework, ColNet, designed to handle both label and task heterogeneity in a decentralized setting.
3 The ColNet Solution
@@
3.1 Problem Statement
In a DFMTL system with $N$ nodes, the learning objective is collaboratively learning task-specific models while preserving data privacy and operating under a fully decentralized communication topology.
@@ 3.1 Problem Statement
3.2 ColNet Architecture
To address the challenges arising from multi-task learning in DFL, the ColNet framework leverages collaborative optimization to enable nodes to benefit from each other while handling diverse tasks. Figure 1 provides an overview of ColNet, which consists of two main phases: a pretraining phase and a DFL learning phase.
@@ 3.2 ColNet Architecture
Pretraining Phase
During the pretraining phase, nodes are clustered according to their data distributions and tasks. Nodes associated with the same task form a single task group. Within each task group, local models adopt the same architecture because they address an identical learning objective. In order to support multi-task learning, all nodes share a unified backbone network, whereas the top (task layers) differ depending on the particular task. Consequently, different task groups utilize a common backbone structure but maintain distinct task layers. Furthermore, each task group designates a task leader node, which serves as a communication bridge in subsequent cross-group information exchanges.
Learning Phase
Once task clustering is complete, the framework proceeds with the DFL learning process, which comprises four iterative steps, illustrated in Figure 1:
@@
    Local Training: Each node trains its local model on the corresponding dataset, updating both the backbone and task-layer parameters.
@@
    Intra-Task Aggregation: After a predefined number of local training epochs, an intra-task aggregation step is triggered. Nodes within the same task group exchange their updated backbone parameters and compute an averaged backbone across the group.
@@
    Cross-Task Aggregation: Each task leader node computes its backbone delta, the difference between the newly averaged backbone and the backbone used at the start of the local training phase, and shares this delta with the leaders of other task groups. Each task leader applies a simplified Hyper Conflict Averse Aggregation (HCA) to the received deltas, merging them into a new backbone.
@@
    Leader Aggregation and Distribution: The updated backbone is then propagated from leader nodes to all members of the corresponding task group for local model updates.
Upon completing these four steps, one training round of ColNet concludes, and the process repeats until a specified termination criterion is met. This iterative procedure allows each node to benefit from both intra-group collaborations with nodes sharing the same task and cross-group exchanges with nodes working on different tasks, thereby enhancing overall learning performance in a decentralized setting. The following subsections detail the steps in each phase.
3.3 Pretraining Phase
This section provides a detailed explanation of the Pretraining Phase in ColNet, covering both the task clustering process and the selection of leader nodes.
Task Clustering
A static grouping strategy is adopted for task clustering, where nodes assigned to the same task are placed in the same group. These nodes share an identical neural network architecture comprising both a backbone network and the task layers. Different task groups use the same backbone structure but employ different task layers to accommodate their respective tasks.
Leader Selection
A rotating-leader approach is adopted to distribute the leader’s computational overhead among group members. One member of each group is randomly selected as the leader node. After each cross-task group aggregation step, the current leader randomly selects a successor and notifies the corresponding client. The new leader then updates all other group leaders, who relay the information to their newly designated leaders.
3.4 Learning Phase
@@
Compared to standard DFL methods, the main novelty of ColNet during training lies in introducing both intra-task and cross-task aggregation. The learning process in each client is presented in Algorithm 1.
Algorithm 1 ColNet Client Learning Process
1:Input: Intra-task peers $K$, leader set $L$, local epochs $epochs$, total rounds $rounds$, boolean $isLeader$
2:for $r=1$ to $rounds$ do
3:     for $e=1$ to $epochs$ do $\triangleright$ Local Training
4:         $\mathbf{w}\leftarrow$ Train locally

### v2

@@ Abstract
The integration of Federated Learning (FL) and Multi-Task Learning (MTL) has been explored to address client heterogeneity, with Federated Multi-Task Learning (FMTL) treating each client as a distinct task. However, most existing research focuses on data heterogeneity (e.g., addressing non-IID data) rather than task heterogeneity, where clients solve fundamentally different tasks. Additionally, much of the work relies on centralized settings with a server managing the federation, leaving the more challenging domain of decentralized FMTL largely unexplored. Thus, this work bridges this gap by proposing ColNet, a framework designed for heterogeneous tasks in decentralized federated environments.
ColNet partitions models into a backbone and task-specific heads, and uses adaptive clustering based on model and data sensitivity to form task-coherent client groups. Backbones are averaged within groups, and group leaders perform hyper-conflict-averse cross-group aggregation. Across datasets and federations, ColNet outperforms competing schemes under label and task heterogeneity and shows robustness to poisoning attacks.
I Introduction
@@
The integration of Federated Learning (FL) and Multi-Task Learning (MTL) has gained significant attention for its ability to tackle diverse but related tasks in a collaborative and privacy-preserving manner. FL trains models across clients by exchanging model updates rather than raw data, preserving data privacy [1]. Decentralized FL (DFL) eliminates the need for a central server, enhancing system robustness by avoiding single points of failure [2]. MTL, traditionally performed on a single machine, leverages task interdependence to improve generalization, reduce overfitting, and address data sparsity [3, 4]. Extending MTL to federated settings results in Federated Multi-Task Learning (FMTL) [5], combining FL’s privacy preservation with MTL’s collaborative task-sharing benefits.
@@
In both centralized and decentralized FMTL, existing research primarily addresses the challenge of non-independent and identically distributed (non-IID) data across clients, a core issue in FL [6]. Non-IID data arises in various forms, including attribute skew (differences in feature distributions), label skew (differences in label distributions), and temporal skew (time-based changes in data distributions), each complicating model training [7]. However, while much progress has been made in handling non-IID data, most studies focus on label heterogeneity rather than deeper task heterogeneity. Many approaches assume that all clients access the same set of class labels, even in non-IID settings. In reality, clients may encounter only subsets of labels relevant to their tasks [8], creating disparities in outputs and limiting collaborative learning. Addressing task heterogeneity, where clients solve fundamentally different problems, remains a significant challenge, especially in decentralized settings.
@@
Aggregation algorithms, like FedAvg, struggle in multi-task learning scenarios due to their reliance on model homogeneity [9]. Task heterogeneity introduces complexities, as clients may require distinct model architectures for tasks (e.g., classification versus point detection), leading to conflicting and dominant gradients during training [10, 11]. While recent methods such as FedBone [12] and FedHCA2 [10] mitigate these challenges, they focus on centralized settings (CFMTL), leaving decentralized FMTL (DFMTL) largely unexplored. Decentralization amplifies these challenges due to the absence of a coordinating server, demanding robust methods for managing tasks and model diversity.
This work proposes ColNet, a decentralized framework for training multi-task models in heterogeneous environments. ColNet partitions models into a backbone and task-specific heads. ColNet uses adaptive clustering that leverages model sensitivity and data sensitivity to automatically partition nodes into task-coherent groups. Within each group, backbone layers are aggregated using FedAvg. Group leaders then exchange the group-averaged backbones across groups through a hyper-conflict-averse (HCA) aggregation scheme, which reduces gradient conflict and strengthens cross-task collaboration.
TABLE I: Overview of Research in Federated Multi-Task Learning
@@
Convex MOCHA [5] Convex optimization, dynamic client inclusion Label heterogeneity
Optimization OFMTL [13] Convex optimization, dynamic client inclusion Label heterogeneity
Regularization Ditto [14] Bi-level optimization, local regularization Label heterogeneity
PMTL [15] Differential privacy, noise-injected global aggregation Label heterogeneity
Privatization FedPer [16] Shared base layers, private personalization layers Label heterogeneity
Centralized FedRep [17] Shared base layers, head freezing Label heterogeneity
Grouping CFL [18] Client grouping, cosine similarity Label heterogeneity
HeurFedAMP [19] Dynamic grouping, cosine similarity Label heterogeneity
Knowledge EFDLS [20] Knowledge distillation, local teacher-student models Label heterogeneity
Distillation FedICT [21] Two-way knowledge distillation, reduced model updates Label heterogeneity
Hybrid FedBone [12] Split learning, gradient projection, task adaptation Task heterogeneity
FedHCA2 [10] Hyper conflict-averse aggregation, encoder-decoder architecture Task heterogeneity
Regularization DCLPMN [22] Graph-based regularization, neighbor collaboration Label heterogeneity
Decentralized Hybrid SpreadGNN [23] Decentralized aggregation, molecular data focus Label heterogeneity
Hybrid ColNet (This work) Adaptive task clustering, shared backbone, hyper conflict-averse aggregation Task heterogeneity
Experiments on the CIFAR-10 and CelebA datasets demonstrate the effectiveness of ColNet. Firstly, the framework’s performance is analyzed with varying aggregation round frequencies, showing that more frequent aggregations improve model performance. Secondly, different levels of layer privatization are examined, revealing that sharing most layers benefits class label heterogeneity setups, while increasing privatization slightly enhances performance in task heterogeneity scenarios. Compared with other schemes, ColNet consistently outperforms baselines in both class label and task heterogeneity scenarios. Additional robustness evaluation indicates that ColNet is resilient to diverse poisoning attacks, including model poisoning and data poisoning, highlighting its capability to address diverse challenges in DFL.
II Related Work
This section provides an overview of FMTL research, covering both centralized and decentralized approaches and comparing them based on the techniques employed and types of heterogeneity addressed, as presented in TABLE I.
@@ A. Centralized Approaches
FMTL was first introduced in 2017 with the MOCHA framework [5], and since then, a variety of approaches have emerged to tackle different aspects of multi-task learning in centralized federated settings. Early work primarily focused on solving convex optimization problems, such as MOCHA and OFMTL [13], which were robust but unsuitable for modern non-convex deep learning methods. Subsequent research addressed this limitation, extending CFMTL to handle non-convex problems and introducing techniques like regularization, privatization, grouping, and knowledge distillation.
Ditto [14] introduced bi-level optimization techniques that allow each client to maintain a personalized model while regularizing it toward a global model. This approach ensures uniform performance across devices and enhances resistance to data and model poisoning attacks. Similarly, PMTL [15] leverages relaxed differential privacy to learn personalized models while safeguarding client data. By adding noise during global aggregation, PMTL enables effective collaboration without compromising privacy. FedPer [16] tackled statistical heterogeneity by dividing models into shared base layers and private personalization layers. Extensions like FedRep [17] introduced mechanisms such as freezing the shared layers during early training phases to improve adaptability and achieve higher test accuracy.
Frameworks like CFL [18] introduced grouping strategies to cluster clients with similar data distributions, reducing the impact of conflicting gradients during aggregation. HeurFedAMP [19] further refined this concept by using cosine similarity to dynamically determine grouping, allowing more granular adaptation to client data heterogeneity. EFDLS [20] utilized knowledge distillation models to transfer knowledge efficiently within individual clients. This method enables collaborative learning by sharing only essential updates, reducing communication overhead. FedICT [21] extended this idea by employing a two-way distillation process that customizes local models for specific tasks while maintaining a flexible and adaptive global model, making it suitable for diverse client environments.
Most existing studies primarily focus on class label heterogeneity, restricting their adaptability in dynamic environments. Frameworks like FedBone [12] introduced split learning to separate general models on a central server from task-specific models on clients. Similarly, FedHCA2 [10] proposed innovative aggregation techniques, such as HCA Aggregation, to address gradient conflicts, but it was designed for centralized environments. These advancements have predominantly relied on centralized coordination, simplifying aggregation but introducing challenges like single points of failure, limited scalability, and privacy risks.
@@ B. Decentralized Approaches
DFMTL eliminates the need for a central server, allowing clients to communicate directly in a peer-to-peer manner. Despite the advantages, DFMTL introduces added challenges, such as managing gradient conflicts, asynchronous updates, and efficient client communication. Early research, such as DCLPMN [22], addressed these challenges by leveraging graph-based regularization to model task relationships, enabling collaboration between neighboring clients. SpreadGNN [23] applied decentralized aggregation techniques tailored for tasks like molecular data analysis to improve scalability and address domain-specific requirements.
As presented in TABLE I, while research on FMTL has made significant progress, most works address label heterogeneity and non-IID data rather than task heterogeneity, where clients solve different tasks. Additionally, decentralized settings remain underexplored compared to centralized frameworks, despite their potential for improved privacy and scalability. Existing frameworks rarely address task heterogeneity in decentralized environments, leaving a clear gap for Heterogeneous DFMTL. This work addresses this gap by proposing a novel framework, ColNet, designed to handle both label and task heterogeneity in a decentralized setting.
III The ColNet Solution
@@
III-A Problem Statement
In a DFMTL system with $N$ nodes, the learning objective is to collaboratively learning task-specific models while preserving data privacy and operating under a fully decentralized communication pattern.
@@ III-A Problem Statement
III-B ColNet Architecture
To tackle the challenges of multi-task learning in DFL, ColNet uses collaborative optimization so that nodes benefit from both task-coherent peers and cross-task exchanges. Figure 1 outlines a five-stage recurrent workflow.
@@ III-B ColNet Architecture
    Local Learning: Each node trains on its local dataset and updates both the shared backbone and its task-specific layers.
@@
    Model Exchanging: Nodes exchange their current backbone parameters with peers in the DFL overlay. Task-specific layers are not shared.
@@
    Adaptive Clustering and Intra-Group Aggregation: Using model-sensitivity and data-sensitivity indicators derived from Stage 2, nodes are automatically grouped into task-coherent clusters. Within each cluster, nodes perform intra-group aggregation on the backbone (FedAvg), while task-specific layers remain private to preserve task diversity. A cluster leader is designated to coordinate cross-group exchange.
@@
    Cross-Group Aggregation: Cluster leaders exchange their group-averaged backbone. Each leader applies a simplified HCA aggregation to reduce conflicting gradients across tasks, producing an updated backbone that integrates complementary information from other groups.
  5. 5.
    Model Redistribution: Leaders redistribute the updated backbone to all members of the corresponding clusters. Nodes synchronize their backbone parameters and resume Stage 1 with refreshed initialization.
Algorithm 1 ColNet Five-Stage Client Workflow (Node ii)
1:Graph $G=(\mathcal{V},\mathcal{E})$, neighbors $\mathcal{N}(i)$, data $\mathcal{D}_{i}$, loss $\ell$, rounds $R$, local epochs $E$, clusters $K$, weights $\alpha,\beta,\gamma$, temps $\tau_{\text{loss}},\tau_{J}$, leader set $\mathcal{L}_{\text{lead}}$
2:Initialize $\mathbf{w}_{i}=(B_{i},H_{i})$; $B_{i}^{\text{prev}}\leftarrow B_{i}$
3:for $r=1$ to $R$ do
4:  Stage 1: Local Learning $\triangleright$ optimize $f_{i}$ on $\mathcal{D}_{i}$
5:  Update $(B_{i},H_{i})$ by $E$-epoch local training with loss $\ell$
6:  Stage 2: Model Exchanging $\triangleright$ backbone only
7:  Exchange $B_{i}$ with all $j\in\mathcal{N}(i)$ to obtain $\{B_{j}\}_{j\in\mathcal{N}(i)}$
8:  Stage 3: Adaptive Clustering & Intra-Group Aggregation
9:  $(g_{i},\,\mathcal{K}(i))\leftarrow\textit{TaskAwareClustering}(B_{i},H_{i},\{B_{j}\},\mathcal{D}_{i})$
10:  $B_{i}\leftarrow\textit{FedAvg}(\{B_{k}\}_{k\in\mathcal{K}(i)}\cup\{B_{i}\})$ $\triangleright$ heads $H$ stay local
11:  Stage 4: Cross-Group Aggregation $\triangleright$ leaders only
12:  if $i\in\mathcal{L}_{\text{lead}}$ then
13:    $\Delta\bar{B}_{i}\leftarrow B_{i}-B_{i}^{\text{prev}}$
14:    $B_{i}\leftarrow\textit{LeaderHCA}(B_{i},\Delta\bar{B}_{i},\mathcal{L}_{\text{lead}})$
15:  end if
16:  Stage 5: Redistribution
17:  if $i\in\mathcal{L}_{\text{lead}}$ then
18:    Broadcast updated $B_{i}$ to all $k\in\mathcal{K}(i)$
19:  else
20:    Receive $B_{\text{leader}}$ from group leader; $B_{i}\leftarrow B_{\text{leader}}$
21:  end if
22:  $B_{i}^{\text{prev}}\leftarrow B_{i}$
23:end for
III-C Learning Process
@@
Compared to standard DFL methods, the main novelty of ColNet during training lies in introducing adaptive clustering, intra-task, and cross-task aggregation. The learning process in each client is presented in Algorithm 1.
Local Learning
Client $i$ solves the local objective on $\mathcal{D}_{i}$ for $E$ epochs (cf. Eq. (1)), updating both $B_{i}$ and $H_{i}$.
Model Exchanging
Clients exchange backbone parameters with neighbors in $G$, i.e., each node broadcasts $B_{i}$ and receives $\{B_{j}\}_{j\in\mathcal{N}(i)}$. Task heads $H$ are not shared.
Adaptive Clustering and Intra-Group Aggregation
Given received backbones $\{B_{j}\}_{j\in\mathcal{N}(i)}$, client $i$ forms temporary models $M_{0}=(B_{i},H_{i})$ and $M_{j}=(B_{j},H_{i})$ for all $j\in\mathcal{N}(i)$ so that all models share the same task head $H_{i}$. This aligns the evaluation to client $i$’s task and avoids confounding from heterogeneous heads. The adaptive clustering is presented in Algorithm 2.
(1) Three signals on local data. On $\mathcal{D}_{i}$, client $i$ computes three pairwise signals among the index set $\mathcal{I}=\{0\}\cup\mathcal{N}(i)$:
Cosine similarity: $C[p,q]\;=\;\frac{\langle\theta(M_{p}),\,\theta(M_{q})\rangle}{\|\theta(M_{p})\|\,\|\theta(M_{q})\|},$ (2)
Empirical loss: $\Lambda[p,q]\;=\;\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}\big[\ell\big(f(x;M_{q}),y\big)\big],$ (3)
Avg. Jacobian norm: $J[p,q]\;=\;\mathbb{E}_{x\sim\mathcal{D}_{i}}\!\left[\left\|\,\frac{\partial f(x;M_{q})}{\partial z(x;M_{q})}\,\right\|_{F}\right].$ (4)
Here $\theta(M)$ stacks the (flattened) backbone parameters of $M$ (heads are identical and omitted), $f$ is the task output (e.g., logits), and $z(\cdot;M_{q})$ denotes the backbone feature right before the head. The Jacobian norm measures the model’s local sensitivity to backbone features, serving as a proxy for model stability.
Together, these three signals capture both model similarity and data-dependent sensitivity on $\mathcal{D}_{i}$, providing a balanced view that is less biased than any single metric.
(2) Symmetrization and temperature scaling. $\Lambda$ and $J$ are asymmetric because they evaluate $M_{q}$ on $\mathcal{D}_{i}$, thus, they are symmetrized via:
$\Lambda\leftarrow\tfrac{1}{2}(\Lambda+\Lambda^{\top})$ (5)
$J\leftarrow\tfrac{1}{2}(J+J^{\top}).$
Then, they are converted into similarities via temperatures $\tau_{\text{loss}},\tau_{J}>0$:
$\Lambda_{\text{sim}}[p,q]=\exp\!\big(-\Lambda[p,q]/\tau_{\text{loss}}\big)$ (6)
$J_{\text{sim}}[p,q]=\exp\!\big(-J[p,q]/\tau_{J}\big).$
In practice, this work standardizes each matrix before exponentiation (z-score per-matrix) to balance scale across signals.
(3) Affinity fusion and self-confidence. This work fuses the three channels into an affinity matrix:
$S\;=\;\alpha\,C\;+\;\beta\,\Lambda_{\text{sim}}\;+\;\gamma\,J_{\text{sim}}$ (7)
$\qquad\alpha,\beta,\gamma\geq 0\;\&\;\alpha{+}\beta{+}\gamma=1,$
(4) Spectral clustering. Apply normalized spectral clustering on the affinity matrix $S$ to obtain $K$ clusters. Let $g_{i}$ be the label assigned to index 0 (client $i$), and define its in-group peers as $\mathcal{K}(i)=\{k\in\mathcal{N}(i)\cup\{i\}\mid g_{k}=g_{i}\}$.
(5) Intra-group aggregation (backbone only). Within $\mathcal{K}(i)$, clients average backbones while keeping heads local:
$\bar{\mathbf{w}}^{B}\;=\;\frac{1}{|\mathcal{K}(i)|}\!\sum_{k\in\mathcal{K}(i)}\mathbf{w}_{k}^{B},\qquad B_{i}\leftarrow\bar{\mathbf{w}}^{B}.$ (8)
This produces a group-coherent backbone $B_{i}$ tailored to $i$’s task via $H_{i}$.
Algorithm 2 TaskAwareClustering at Node $i$
[…]
2:Cluster label $g_{i}$, peer set $\mathcal{K}(i)$
3:Build models $M_{0}\!\leftarrow\!(B_{i},H_{i})$ and $M_{j}\!\leftarrow\!(B_{j},H_{i})$ for $j\!\in\!\mathcal{N}(i)$
4:Let index set $\mathcal{I}\!\leftarrow\!\{0\}\cup\mathcal{N}(i)$; init $C,\Lambda,J\in\mathbb{R}^{|\mathcal{I}|\times|\mathcal{I}|}$
5:for all $p,q\in\mathcal{I}$ do
6:  $C[p,q]\leftarrow\textit{CosSim}(M_{p},M_{q})$
7:  $\Lambda[p,q]\leftarrow\textit{AvgLoss}(M_{q};\mathcal{D}_{i},\ell)$
8:  $J[p,q]\leftarrow\textit{AvgJacobianNorm}(M_{q};\mathcal{D}_{i})$
9:end for
[…]
14:Apply spectral clustering on $S$ into $K$ clusters; obtain $g_{i}$
[…]
16:return $(g_{i},\,\mathcal{K}(i))$
(6) Leader Selection. Once task groups are established, each group appoints a leader on a rotating basis to balance coordination overhead. In round $r$, one member serves as leader and coordinates inter-group aggregation; after this step, leadership rotates (round-robin) to the next member in the group. The outgoing leader announces the handover to its group and to peer leaders so that subsequent cross-group exchanges use the updated routing.
Cross-Group Aggregation
Cross-group aggregation is challenging due to potential conflicts between updates produced by heterogeneous tasks. Let $\mathcal{L}_{\text{lead}}$ denote the set of current group leaders and $G_{\ell}=|\mathcal{L}_{\text{lead}}|$ the number of groups. Each leader $i\!\in\!\mathcal{L}_{\text{lead}}$ computes a backbone delta to its previous round,
$\Delta\mathbf{w}_{i}^{B}\;=\;\bar{\mathbf{w}}_{i}^{B}-\mathbf{w}_{i,\text{prev}}^{B},$ (9)
exchanges $\Delta\mathbf{w}_{i}^{B}$ with other leaders, and forms the mean update $\Delta\bar{\mathbf{w}}^{B}=\tfrac{1}{G_{\ell}}\sum_{\ell\in\mathcal{L}_{\text{lead}}}\Delta\mathbf{w}_{\ell}^{B}.$
Inspired by FedHCA2 [10], ColNet adapts a gradient-alignment objective that seeks an aggregated update $\tilde{U}$ close to $\Delta\bar{\mathbf{w}}^{B}$ while improving its alignment with the worst-case task direction:
$\text{s.t.}\quad\|\tilde{U}-\Delta\bar{\mathbf{w}}^{B}\|\;\leq\;c\,\|\Delta\bar{\mathbf{w}}^{B}\|$
where $c\in[0,1)$ bounds the deviation from the mean update. A Lagrangian treatment yields the closed-form
$\tilde{U}\;=\;\Delta\bar{\mathbf{w}}^{B}\;+\;\frac{\sqrt{\phi}}{\|U_{w}\|}\,U_{w}$ (11)
$U_{w}\;=\;\frac{1}{G_{\ell}}\sum_{\ell\in\mathcal{L}_{\text{lead}}}w_{\ell}\,\Delta\mathbf{w}_{\ell}^{B}$
$\phi\!=\!c^{2}\|\Delta\bar{\mathbf{w}}^{B}\|^{2}$
with nonnegative weights $w_{\ell}$ determined by the chosen conflict metric. This update balances proximity to the mean and reduction of cross-task conflicts, as shown in Algorithm 3. Each leader applies $\tilde{U}$ to its backbone and proceeds to redistribution.
Algorithm 3 LeaderHCA (Cross-Group Agg. at Leader ii)
1:Current backbone $B_{i}$, local delta $\Delta\bar{B}_{i}$, leader set $\mathcal{L}_{\text{lead}}$
2:Updated backbone $B_{i}$
3:for all $\ell\in\mathcal{L}_{\text{lead}}\setminus\{i\}$ do
4:  Send $\Delta\bar{B}_{i}$ to $\ell$; Receive $\Delta\bar{B}_{\ell}$ from $\ell$